Suppose that you invested $100,000 in a business venture a number of years ago, your annual rate of return has been 11% and you have been given an offer of $283,942.00 to purchase your share of the business.  How many years (round to the nearest whole year) have you been invested in the business (i.e. assume you invested on January 1, 200X and it is now December 31, 20XX)?

Group of answer choices

A. 8 years

D. 10 years

B. 11 years

C. 9 years
Answer #1

Future Value = Present Value * (1+(Rate per year/no. of compounding per year)^(no. of years*no. of compounding per year)

283942 = 100000*1.11^n

1.11^n = 283942/100000

= 2.83942

n = log2.83942 / log1.11

= 0.45322963704/0.04532297878

= 9.99999667365

=  10 years